OpenStudy (anonymous):

In how many ways can a committee of 5 people be chosen from a total of 30 possible people? 105, 729 112,098 123,567 142,506

OpenStudy (cwrw238):

30 C 5 = 30*29*28*27*26 --------------- 5*4*3*2*1

OpenStudy (lochana):

30!/(25!)(5!)=(30*29*28*27*26)/(5*4*3*2*1)

OpenStudy (lochana):

142,506

OpenStudy (nurali):

Ways=30C5=30*29*28*27*26/5*4*3*2*1

OpenStudy (wolf1728):

combinations = 30! / (5!) * (25!) combinations = 30*29*28*27*26/ 5*4*3*2*1 combinations = 30*29*28*27*26 / 120 =17100720/120 =142,506